Diffusion Morning Report 08/09/07

Ted Berg tberg at stanford.edu
Thu Aug 9 06:28:41 PDT 2007


Good Morning All,
    Looking OK this morning. POCl3 and BBr3 under observation. MFC 
changed on TEOS2 checking operation. Poly still has pressure control 
issues.. No other issues at this time.

  	  	  	  	  	 
  	Date: 	8/9/07 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	  	  	  	 
Tylan 1 	up 	  	  	  	 
Tylan 2 	up 	  	  	  	 
Tylan 3 	up 	  	  	  	 
Tylan 4 	up 	  	  	  	 
Tylan 5-BBr3 	yellow 	observation 	  	  	 
Tylan 6-POCl3 	yellow 	observation 	  	  	 
Tylan 7-TEOS2 	yellow 	mfc changed needs test 	9.99/30 	  	 
Tylan 8-SiGe 	up 	  	19.63/30 	  	 
Tylan 9-Poly 	down 	N2 purge issue 	17.2/30 	  	 
Tylan 10-Nitride 	up 	  	26.44/30 	  	 
Tylan 11-BPSG-LTO 	up 	  	1.95/30 	  	 
Tylan 12-FGA 	up 	  	  	  	 
  	  	  	  	  	 
Thermco 1 	up 	  	  	  	 
Thermco 2 	up 	  	  	  	 
Thermco 3 	future 	  	  	  	 
Thermco4 	future 	  	  	  	 
Thermco Poly1 	up 	  	  	  	 
Thermco Poly2 	future 	  	
	
	
Thermco Nitride1 	up 	  	  	  	 
Thermco Nitride2 	future 	
	  	  	 
Burn Box 	Up 	  	   	  	 
Ebeam 	up 	  	  	  	 
ZYGO 	up 	  	  	  	 
FGA2 	up 	  	  	  	 
SiH4 for furnaces 	up 	A on line 580 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20070809/6e1d9541/attachment.html>


More information about the diffusion mailing list